OpenStudy (anonymous):

Solve for t : (1/2)(-32t^2+t+400)

sam (.sam.):

no you cant

OpenStudy (anonymous):

How?

sam (.sam.):

you mean simplify?

sam (.sam.):

if its solve then (1/2)(-32t^2+t+400)=0

sam (.sam.):

\[\begin{array}{l} \frac{1}{2} \left(-32 t^2+t+400\right)=0 \\ \text{Divide both sides by }\frac{1}{2} \\ -32 t^2+t+400=0 \\ \text{Use quadratic formula \to solve the equation:} x=\frac{-b \pm \sqrt{b^2-4ac}}{2a} \\\text{Where a=-32,b=1,c=400}\\\end{array}\]

sam (.sam.):

\[\huge x=\frac{-b \pm \sqrt{b^2-4ac}}{2a}\] \[\huge x=\frac{-1 \pm \sqrt{1^2-4(-32)(400)}}{2(-32)}\] \[\huge x=\frac{1 \pm \sqrt{51201}}{64}\] \[\huge x=\frac{1 + \sqrt{51201}}{64}~~~or~~~\huge x=\frac{1 - \sqrt{51201}}{64}\]

sam (.sam.):

just replace x as t, its the same

OpenStudy (anonymous):

Thank you